ROTATIONALLY RESOLVED STUDIES OF P ARA-AMINOPHENOL AND ITS COMPLEXES.^{a}

JUSTIN W. YOUNG, \underline{PATRICK S. WALSH}, ADAM J. FLEISHER and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pa 15260.

In biological systems, solvent molecules play a significant role through non-covalent interactions. From this, the desire to understand the properties of solvent molecules bound to different functional groups of a bio-molecule follows. The goal of this study is to use rotationally resolved electronic spectroscopy to determine the position of attachment and intermolecular motions of a single solvent molecule linked to three possible receptor sites in para-aminophenol; the hydroxyl group, the amino group, and the aromatic ring.

aWork supported by NSF(CHE-0911117)